OpenStudy (anonymous):

what is the solution of the system of equations? 2x + 2y + 3z = -6 3x + 5y + 4z = 3 2x + 3y +4z = -10

OpenStudy (anonymous):

Subtract the first equation with the last equation like this: \[(2x+2y+3z)-(2x+3y+4z)=(-6)-(-10)\]

OpenStudy (anonymous):

ohh ok. I can't believe I didn't notice that. I was so focused on the first and second one.

OpenStudy (anonymous):

We're not done yet.

OpenStudy (anonymous):

You have to multiply the first equation by 3 and then you would have to multiply the second equation by 2. \[3(2x+2y+3z)=3(-6)\] \[2(3x+5y+4z)=2(3)\]

OpenStudy (anonymous):

Then you subtract to get rid of the x. Then you would have two equations with two variables: y and z.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then you can go on from there to solve those two new equations like a system.

OpenStudy (anonymous):

ok thanks so much for explaining.

OpenStudy (anonymous):

No worries.